# A lower bound for the Weisfeiler-Leman dimension of circulant graphs

### v1

@@ 9.1. The main theorem
In this section, we construct an S-ring ${\cal A}^{\star}$ and a set of ${\cal A}^{\star}$-multipliers. They will be used to prove Theorem 1.1 in Section 10. The relevant properties of ${\cal A}^{\star}$ are given in the theorem below.
@@
Let ${\cal G}\in{\mathfrak{B}}_{a}$ and $k$ a positive integer. Assume that each separator of the graph ${\cal G}$ has cardinality greater than $k$. Then there exist
@@
    a totally normal circulant S-ring ${\cal A}^{\star}$ of degree $n$ defined as in Theorem 1.1, such that the scheme ${\cal X}(({\cal A}^{\star})_{0})$ is binary,
@@
    an outer ${\cal A}^{\star}$-multiplier ${\mathfrak{c}}$ such that $\operatorname{Iso}({\cal A}^{\star},\varphi_{\mathfrak{c}})=\varnothing$, and
@@
    a $\lfloor k/6\rfloor$-local system of inner ${\cal A}^{\star}$-multipliers with respect to ${\mathfrak{c}}$.
@@
9.2. Construction of the S-ring ${\cal A}^{\star}$
Let ${\cal X}$ be the coherent configuration defined in Subsection 8.1. By the condition (K3), we have $\operatorname{Aut}({\cal X})\leq K_{V}$. Let
${\mathfrak{M}}=\mu(\operatorname{Aut}({\cal X})),$
@@ 9.2. Construction of the S-ring ${\cal A}^{\star}$
${\mathfrak{M}}_{v,v^{\prime}}=\begin{cases}\{(x,y)\in K_{v}\times K_{v^{\prime}}:\ \rho_{v}(x)=\rho_{v^{\prime}}(y)\}&\text{if $(v,v^{\prime})\in E$,}\\ K_{v}\times K_{v^{\prime}}&\text{otherwise.}\\ \end{cases}$
Let ${\cal A}$ be the S-ring defined in Subsection 8.2. Denote by ${\cal A}^{\star}$ the algebraic fusion of ${\cal A}$ with respect to the subgroup ${\mathfrak{M}}\leq{\mathfrak{M}}({\cal A})$,
${\cal A}^{\star}={\cal A}^{\mathfrak{M}}.$
Since ${\cal A}$ and ${\cal A}^{\star}$ have the same ${\cal A}$-subgroups (formula (25)), the subsets $D_{i,j}$ defined by formula (45) are ${\cal A}^{\star}$-sets for all $i,j$.
@@
(48) ${\mathfrak{S}}_{0}({\cal A}^{\star})={\mathfrak{S}}_{0}({\cal A})\quad\text{and}\quad({\cal A}^{\star})_{0}={\cal A}.$
@@
The S-ring ${\cal A}^{\star}$ is totally normal.
@@
Let ${\mathfrak{s}}\in{\mathfrak{S}}_{0}({\cal A}^{\star})$. Then ${\mathfrak{s}}\in{\mathfrak{S}}_{0}({\cal A})$ by the left-hand side of formula (48). By Lemma 8.3(2), this implies that ${\mathfrak{s}}\preceq{\mathfrak{s}}_{u,w}$ for some $u\in V_{1}$ and $w\in V_{2}$. Since by the above, the S-ring $({\cal A}^{\star})_{\mathfrak{s}}=\operatorname{cyc}({\mathfrak{M}}_{\mathfrak{s}},{\mathfrak{s}})$ is normal, we are done. ∎
@@ 9.3. An outer multiplier
In the notation of Subsection 8.1, fix an arbitrary edge $e_{0}=\{u_{i_{0}},w_{j_{0}}\}$ of the graph ${\cal G}$ and the algebraic automorphism $\psi=\psi_{e_{0}}$ of the coherent configuration ${\cal X}$. From condition (K4), it follows that the group $\operatorname{Aut}({\cal X})^{\Omega_{e_{0}}}$ has index $2$ in the group $\bar{K}_{u_{i_{0}}}\times\bar{K}_{w_{j_{0}}}$. Choose any permutation $\bar{k}_{0}$ of this group, not belonging to $\operatorname{Aut}({\cal X})^{\Omega_{e_{0}}}$, and define the permutation $k_{0}\in\operatorname{Aut}({\mathfrak{s}}_{e_{0}})$ by the conditions
@@
Let $u\in V_{1}$ and $w\in V_{2}$. Then the section ${\mathfrak{s}}={\mathfrak{s}}_{u,w}$ belongs to ${\mathfrak{S}}_{0}({\cal A})$, see Lemma 8.3(2). By formula (48), this implies that ${\mathfrak{s}}\in{\mathfrak{S}}_{0}({\cal A}^{\star})$. Let us define a coset ${\mathfrak{c}}_{\mathfrak{s}}$ of the subgroup ${\mathfrak{M}}_{\mathfrak{s}}\leq\operatorname{Aut}({\mathfrak{s}})$ by the formula
(49) ${\mathfrak{c}}_{\mathfrak{s}}=\begin{cases}k_{0}\cdot{\mathfrak{M}}_{\mathfrak{s}}&\text{if ${\mathfrak{s}}={\mathfrak{s}}_{e_{0}}$,}\\ {\mathfrak{M}}_{\mathfrak{s}}&\text{otherwise,}\\ \end{cases}$
where ${\mathfrak{s}}_{e_{0}}={\mathfrak{s}}_{u_{i_{0}},w_{j_{0}}}$. We extend this definition to all sections ${\mathfrak{s}}\in{\mathfrak{S}}_{0}({\cal A}^{\star})$ by setting ${\mathfrak{c}}_{\mathfrak{s}}={\mathfrak{c}}_{{\mathfrak{s}}_{u,w}}$ for suitable $u,w$ (see Lemma 8.3(2) and formula (48)). Furthermore, by formula (24), we have
$({\cal A}^{\star})_{\mathfrak{s}}=\operatorname{cyc}({\mathfrak{M}}_{\mathfrak{s}},{\mathfrak{s}}).$
This S-ring is normal by Lemma 9.3 and hence ${\mathfrak{M}}_{\mathfrak{s}}=\operatorname{Aut}_{{\cal A}^{\star}}({\mathfrak{s}})$ by Lemma 5.2(3). Using arguments in the proof of Lemma 8.4, one can verify that the family ${\mathfrak{c}}:={\mathfrak{c}}(e_{0})$ of the cosets ${\mathfrak{c}}_{\mathfrak{s}}$, ${\mathfrak{s}}\in{\mathfrak{S}}_{0}({\cal A}^{\star})$, is an outer ${\cal A}^{\star}$-multiplier.
In the following statement, we establish a relationship between the isomorphisms of the coherent configuration ${\cal X}$, the corresponding inner multipliers in ${\mathfrak{M}}$, and the outer multiplier ${\mathfrak{c}}$.
@@
(50) $f^{\Omega_{e}}\in\operatorname{Iso}({\cal X}_{\Omega_{e}},\psi_{\Omega_{e}})\quad\Leftrightarrow\quad\text{either }e\not\in E\text{ or }({\mathfrak{m}}_{f})_{e}\in{\mathfrak{c}}_{e},$
@@
We note that if $e\not\in E$, then $\operatorname{Iso}({\cal X}_{\Omega_{e}},\psi_{\Omega_{e}})=\bar{K}_{v}\times\bar{K}_{v^{\prime}}$ and $f^{\Omega_{e}}\in\operatorname{Iso}({\cal X}_{\Omega_{e}},\psi_{\Omega_{e}})$ by condition (K5). Assume that $e\in E$.
Let $e\neq e_{0}$. Then $\operatorname{Iso}({\cal X}_{\Omega_{e}},\psi_{\Omega_{e}})=\operatorname{Aut}({\cal X}_{\Omega_{e}})$ by condition (K4). Therefore, $f^{\Omega_{e}}\in\operatorname{Iso}({\cal X}_{\Omega_{e}},\psi_{\Omega_{e}})$ if and only if
$({\mathfrak{m}}_{f})_{e}\in\mu_{e}(\operatorname{Aut}({\cal X}_{\Omega_{e}}))={\mathfrak{M}}_{e}={\mathfrak{M}}_{{\mathfrak{s}}_{v,v^{\prime}}}={\mathfrak{c}}_{{\mathfrak{s}}_{v,v^{\prime}}}={\mathfrak{c}}_{e},$
@@
Finally, let $e=e_{0}$, i.e., $\{v,v^{\prime}\}=\{u_{i_{0}},w_{j_{0}}\}$. Then by condition (K4), the group $\operatorname{Aut}({\cal X}_{\Omega_{e}})$ has index $2$ in the group $\bar{K}_{u_{i_{0}}}\times\bar{K}_{w_{j_{0}}}$. By virtue of condition (K5), we have
$\operatorname{Iso}({\cal X}_{\Omega_{e}},\psi_{\Omega_{e}})=\bar{k}_{0}\cdot\operatorname{Aut}({\cal X}_{\Omega_{e}}),$
where $\bar{k}_{0}$ is the permutation chosen at the beginning of this subsection. Thus, $f^{\Omega_{e}}\in\operatorname{Iso}({\cal X}_{\Omega_{e}},\psi_{\Omega_{e}})$ if and only if
$({\mathfrak{m}}_{f})_{e}\in\mu_{e}(\bar{k}_{0}\cdot\operatorname{Aut}({\cal X}_{\Omega_{e_{0}}}))={\mathfrak{M}}_{e_{0}}={\mathfrak{M}}_{{\mathfrak{s}}_{u_{i_{0}},w_{j_{0}}}}={\mathfrak{c}}_{{\mathfrak{s}}_{u_{i_{0}},w_{j_{0}}}}={\mathfrak{c}}_{e_{0}},$
@@
Let ${\cal A}^{\star}$ be the S-ring over the group $G$, defined in Subsection 9.2. By Lemma 9.3, ${\cal A}^{\star}$ is totally normal. By formula (48), we have  $({\cal A}^{\star})_{0}={\cal A}$. Hence the scheme ${\cal X}(({\cal A}^{\star})_{0})$ is binary by Lemma 8.2. Thus it suffices to verify that statements (2) and (3) of the theorem in question hold true for the outer ${\cal A}^{\star}$-multiplier ${\mathfrak{c}}$ defined in Subsection 9.3.
To prove statement (2), assume on the contrary that there is $g\in\operatorname{Iso}({\cal A}^{\star},\varphi_{\mathfrak{c}})$. Without loss of generality, we may assume that $g$ is a normalized isomorphism. Since $G$ is of odd order and ${\cal A}^{\star}$ is totally normal, the family
${\mathfrak{m}}=\{g^{\mathfrak{s}}:\ {\mathfrak{s}}\in{\mathfrak{S}}_{0}({\cal A}^{\star})\}$
is an inner ${\cal A}^{\star}$-multiplier, see the end of Subsection 6.2. Furthermore, ${\mathfrak{m}}$ is an inner ${\cal A}$-multiplier by the first equality in (48). In addition, ${\mathfrak{m}}\leq{\mathfrak{c}}$ and hence
@@
By Lemma 8.5, this shows that $f=\mu^{-1}({\mathfrak{m}})$ is an isomorphism of the coherent configuration ${\cal X}$, belonging to $K_{V}$. Besides, let $e=\{v,v^{\prime}\}$, where $v,v^{\prime}\in V$ and $v\neq v^{\prime}$. Then
@@
where $\pi_{v}$ and $\pi_{v^{\prime}}$ are the isomorphisms defined in formula (47). By Lemma 9.4, this implies that $f^{\Omega_{e}}\in\operatorname{Iso}({\cal X}_{\Omega_{e}},\psi_{\Omega_{e}})$. It follows that $f\in\operatorname{Iso}({\cal X},\psi)$, which contradicts Lemma 8.1.
Let us prove statement (3). Put $m=\lfloor k/6\rfloor$. The graph ${\cal G}$ has no separators of cardinality at most $3\cdot(2m)\leq k$. By Lemma 8.1, there exists an $(2m)$-local system
(51) $\{f_{\Delta}\in K_{V}:\Delta\in F_{2m}({\cal X})\}$
of isomorphisms for the coherent configuration ${\cal X}$ with respect to the algebraic isomorphism $\psi$.
Let ${\mathfrak{S}}\subseteq{\mathfrak{S}}_{m}$, where ${\mathfrak{S}}_{m}={\mathfrak{S}}_{m}({\cal A}^{\star})$. By formula in (48) and Lemma 8.3(2), there is a minimal (by inclusion) subset $\bar{V}=\bar{V}({\mathfrak{S}})$ of the set $V$ such that if ${\mathfrak{s}}\in{\mathfrak{S}}$, then ${\mathfrak{s}}\preceq{\mathfrak{s}}_{u,w}$ for some $u\in V_{1}\cap\bar{V}$ and $w\in V_{2}\cap\bar{V}$. Clearly,
@@
and hence $\Omega_{\bar{V}}\in F_{2m}({\cal X})$. Thus the $(2m)$-local system (51) contains an isomorphism $f_{\Delta}\in K_{V}$ of the coherent configuration ${\cal X}$, where $\Delta=\Omega_{\bar{V}}$. Denote by ${\mathfrak{m}}({\mathfrak{S}})$ the inner ${\cal A}$-multiplier $\mu(f_{\Delta})$.
@@
$\{{\mathfrak{m}}({\mathfrak{S}})\in{\mathfrak{M}}({\cal A}^{\star}):\ {\mathfrak{S}}\in{\mathfrak{S}}_{m}\}$
is an $m$-local system of inner ${\cal A}^{\star}$-multipliers with respect to the outer ${\cal A}^{\star}$-multiplier ${\mathfrak{c}}$.
To verify the first relation in (28), let ${\mathfrak{S}}\in{\mathfrak{S}}_{m}$ and ${\mathfrak{s}}\in{\mathfrak{S}}$. By Lemma 8.3(2), we may assume without loss of generality that ${\mathfrak{s}}={\mathfrak{s}}_{u,w}$ for some vertices $u\in V_{1}\cap\bar{V}$ and $w\in V_{2}\cap\bar{V}$, where $\bar{V}=\bar{V}({\mathfrak{S}})$. Since $u,w\in\bar{V}$, we have $\Omega_{u,w}\subseteq\Omega_{\bar{V}}:=\Delta$. By the definition of $f_{\Delta}$, this implies that $(f_{\Delta})^{\Omega_{e}}\in\operatorname{Iso}({\cal X}_{\Omega_{e}},\psi_{\Omega_{e}})$, where $e=\{u,w\}$. By Lemma 9.4, either $e\not\in E$, or $e\in E$ and
@@
To verify the second relation in (28), let ${\mathfrak{S}},{\mathfrak{S}}^{\prime}\in{\mathfrak{S}}_{m}$. Put $\Delta=\Omega_{\bar{V}}$ and $\Delta^{\prime}=\Omega_{\bar{V}^{\prime}}$, where $\bar{V}=\bar{V}({\mathfrak{S}})$ and $\bar{V}^{\prime}=\bar{V}({\mathfrak{S}}^{\prime})$. As above, we have $\Delta,\Delta^{\prime}\in F_{2m}({\cal X})$. By the definition of $(2m)$-local system, this implies that
@@
for some $h_{\Delta,\Delta^{\prime}}\in\operatorname{Aut}({\cal X})$. It follows that $f_{\Delta}{f_{\Delta^{\prime}}}^{-1}=f_{\Delta,\Delta^{\prime}}h_{\Delta,\Delta^{\prime}}$, where the isomorphism $f_{\Delta,\Delta^{\prime}}:=f_{\Delta}{f_{\Delta^{\prime}}}^{-1}h_{\Delta,\Delta^{\prime}}^{-1}$ belongs to $K_{V}$ and is identical on $\Delta\cap\Delta^{\prime}$. Applying the monomorphism $\mu$ from Lemma 8.5 to the last equality, we obtain
@@
By Lemma 6.2, the inner ${\cal A}^{\star}$-multiplier ${\mathfrak{m}}({\mathfrak{S}},{\mathfrak{S}}^{\prime}):=\mu(h_{\Delta,\Delta^{\prime}})$ belongs to the group ${\mathfrak{M}}\leq{\mathfrak{M}}_{aut}({\cal A}^{\star})$. Moreover, the inner ${\cal A}^{\star}$-multiplier $\mu(f_{\Delta,\Delta^{\prime}})$ is identical on ${\mathfrak{S}}\cap{\mathfrak{S}}^{\prime}$. Thus, taking restriction of the equality (52) to ${\mathfrak{S}}\cap{\mathfrak{S}}^{\prime}$, we get

### v2

@@ 9.1. The main theorem
[…]
@@
Let ${\mathcal{G}}\in{\mathfrak{B}}_{a}$ and $k$ a positive integer. Assume that each separator of the graph $\mathcal{G}$} has cardinality greater than $k$. Then there exist
@@
    a totally normal circulant S-ring $\mathcal{A}$$^{\star}$ of degree $n$ defined as in Theorem 1.1, such that the scheme ${\mathcal{X}}(({\mathcal{A}}^{\star})_{0})$ is binary,
@@
[…]
@@
[…]
@@
9.2. Construction of the S-ring $\mathcal{A}$$^{\star}$
Let $\mathcal{X}$} be the coherent configuration defined in Subsection 8.1. By the condition (K3), we have $\operatorname{Aut}({\mathcal{X}})\leq K_{V}$. Let
${\mathfrak{M}}=\mu(\operatorname{Aut}({\mathcal{X}})),$
@@ 9.2. Construction of the S-ring $\mathcal{A}$$^{\star}$
[…]
Let $\mathcal{A}$} be the S-ring defined in Subsection 8.2. Denote by $\mathcal{A}$$^{\star}$ the algebraic fusion of $\mathcal{A}$} with respect to the subgroup ${\mathfrak{M}}\leq{\mathfrak{M}}({\mathcal{A}})$,
${\mathcal{A}}^{\star}={\mathcal{A}}^{\mathfrak{M}}.$
[…]
@@
(48) ${\mathfrak{S}}_{0}({\mathcal{A}}^{\star})={\mathfrak{S}}_{0}({\mathcal{A}})\quad\text{and}\quad({\mathcal{A}}^{\star})_{0}={\mathcal{A}}.$
@@
The S-ring $\mathcal{A}$$^{\star}$ is totally normal.
@@
Let ${\mathfrak{s}}\in{\mathfrak{S}}_{0}({\mathcal{A}}^{\star})$. Then ${\mathfrak{s}}\in{\mathfrak{S}}_{0}({\mathcal{A}})$ by the left-hand side of formula (48). By Lemma 8.3(2), this implies that ${\mathfrak{s}}\preceq{\mathfrak{s}}_{u,w}$ for some $u\in V_{1}$ and $w\in V_{2}$. Since by the above, the S-ring $({\mathcal{A}}^{\star})_{\mathfrak{s}}=\operatorname{cyc}({\mathfrak{M}}_{\mathfrak{s}},{\mathfrak{s}})$ is normal, we are done. ∎
@@ 9.3. An outer multiplier
In the notation of Subsection 8.1, fix an arbitrary edge $e_{0}=\{u_{i_{0}},w_{j_{0}}\}$ of the graph $\mathcal{G}$} and the algebraic automorphism $\psi=\psi_{e_{0}}$ of the coherent configuration $\mathcal{X}$}. From condition (K4), it follows that the group $\operatorname{Aut}({\mathcal{X}})^{\Omega_{e_{0}}}$ has index $2$ in the group $\bar{K}_{u_{i_{0}}}\times\bar{K}_{w_{j_{0}}}$. Choose any permutation $\bar{k}_{0}$ of this group, not belonging to $\operatorname{Aut}({\mathcal{X}})^{\Omega_{e_{0}}}$, and define the permutation $k_{0}\in\operatorname{Aut}({\mathfrak{s}}_{e_{0}})$ by the conditions
@@
Let $u\in V_{1}$ and $w\in V_{2}$. Then the section ${\mathfrak{s}}={\mathfrak{s}}_{u,w}$ belongs to ${\mathfrak{S}}_{0}({\mathcal{A}})$, see Lemma 8.3(2). By formula (48), this implies that ${\mathfrak{s}}\in{\mathfrak{S}}_{0}({\mathcal{A}}^{\star})$. Let us define a coset ${\mathfrak{c}}_{\mathfrak{s}}$ of the subgroup ${\mathfrak{M}}_{\mathfrak{s}}\leq\operatorname{Aut}({\mathfrak{s}})$ by the formula
[…]
where ${\mathfrak{s}}_{e_{0}}={\mathfrak{s}}_{u_{i_{0}},w_{j_{0}}}$. We extend this definition to all sections ${\mathfrak{s}}\in{\mathfrak{S}}_{0}({\mathcal{A}}^{\star})$ by setting ${\mathfrak{c}}_{\mathfrak{s}}={\mathfrak{c}}_{{\mathfrak{s}}_{u,w}}$ for suitable $u,w$ (see Lemma 8.3(2) and formula (48)). Furthermore, by formula (24), we have
$({\mathcal{A}}^{\star})_{\mathfrak{s}}=\operatorname{cyc}({\mathfrak{M}}_{\mathfrak{s}},{\mathfrak{s}}).$
[…]
In the following statement, we establish a relationship between the isomorphisms of the coherent configuration $\mathcal{X}$}, the corresponding inner multipliers in $\mathfrak{M}$}, and the outer multiplier $\mathfrak{c}$}.
@@
(50) $f^{\Omega_{e}}\in\operatorname{Iso}({\mathcal{X}}_{\Omega_{e}},\psi_{\Omega_{e}})\quad\Leftrightarrow\quad\text{either }e\not\in E\text{ or }({\mathfrak{m}}_{f})_{e}\in{\mathfrak{c}}_{e},$
@@
We note that if $e\not\in E$, then $\operatorname{Iso}({\mathcal{X}}_{\Omega_{e}},\psi_{\Omega_{e}})=\bar{K}_{v}\times\bar{K}_{v^{\prime}}$ and $f^{\Omega_{e}}\in\operatorname{Iso}({\mathcal{X}}_{\Omega_{e}},\psi_{\Omega_{e}})$ by condition (K5). Assume that $e\in E$.
Let $e\neq e_{0}$. Then $\operatorname{Iso}({\mathcal{X}}_{\Omega_{e}},\psi_{\Omega_{e}})=\operatorname{Aut}({\mathcal{X}}_{\Omega_{e}})$ by condition (K4). Therefore, $f^{\Omega_{e}}\in\operatorname{Iso}({\mathcal{X}}_{\Omega_{e}},\psi_{\Omega_{e}})$ if and only if
$({\mathfrak{m}}_{f})_{e}\in\mu_{e}(\operatorname{Aut}({\mathcal{X}}_{\Omega_{e}}))={\mathfrak{M}}_{e}={\mathfrak{M}}_{{\mathfrak{s}}_{v,v^{\prime}}}={\mathfrak{c}}_{{\mathfrak{s}}_{v,v^{\prime}}}={\mathfrak{c}}_{e},$
@@
Finally, let $e=e_{0}$, i.e., $\{v,v^{\prime}\}=\{u_{i_{0}},w_{j_{0}}\}$. Then by condition (K4), the group $\operatorname{Aut}({\mathcal{X}}_{\Omega_{e}})$ has index $2$ in the group $\bar{K}_{u_{i_{0}}}\times\bar{K}_{w_{j_{0}}}$. By virtue of condition (K5), we have
$\operatorname{Iso}({\mathcal{X}}_{\Omega_{e}},\psi_{\Omega_{e}})=\bar{k}_{0}\cdot\operatorname{Aut}({\mathcal{X}}_{\Omega_{e}}),$
where $\bar{k}_{0}$ is the permutation chosen at the beginning of this subsection. Thus, $f^{\Omega_{e}}\in\operatorname{Iso}({\mathcal{X}}_{\Omega_{e}},\psi_{\Omega_{e}})$ if and only if
$({\mathfrak{m}}_{f})_{e}\in\mu_{e}(\bar{k}_{0}\cdot\operatorname{Aut}({\mathcal{X}}_{\Omega_{e_{0}}}))={\mathfrak{M}}_{e_{0}}={\mathfrak{M}}_{{\mathfrak{s}}_{u_{i_{0}},w_{j_{0}}}}={\mathfrak{c}}_{{\mathfrak{s}}_{u_{i_{0}},w_{j_{0}}}}={\mathfrak{c}}_{e_{0}},$
@@
[…]
To prove statement (2), assume on the contrary that there is $g\in\operatorname{Iso}({\mathcal{A}}^{\star},\varphi_{\mathfrak{c}})$. Without loss of generality, we may assume that $g$ is a normalized isomorphism. Since $G$ is of odd order and $\mathcal{A}$$^{\star}$ is totally normal, the family
${\mathfrak{m}}=\{g^{\mathfrak{s}}:\ {\mathfrak{s}}\in{\mathfrak{S}}_{0}({\mathcal{A}}^{\star})\}$
[…]
@@
By Lemma 8.5, this shows that $f=\mu^{-1}({\mathfrak{m}})$ is an isomorphism of the coherent configuration $\mathcal{X}$}, belonging to $K_{V}$. Besides, let $e=\{v,v^{\prime}\}$, where $v,v^{\prime}\in V$ and $v\neq v^{\prime}$. Then
@@
where $\pi_{v}$ and $\pi_{v^{\prime}}$ are the isomorphisms defined in formula (47). By Lemma 9.4, this implies that $f^{\Omega_{e}}\in\operatorname{Iso}({\mathcal{X}}_{\Omega_{e}},\psi_{\Omega_{e}})$. It follows that $f\in\operatorname{Iso}({\mathcal{X}},\psi)$, which contradicts Lemma 8.1.
Let us prove statement (3). Put $m=\lfloor k/6\rfloor$. The graph $\mathcal{G}$} has no separators of cardinality at most $3\cdot(2m)\leq k$. By Lemma 8.1, there exists an $(2m)$-local system
(51) $\{f_{\Delta}\in K_{V}:\Delta\in F_{2m}({\mathcal{X}})\}$
of isomorphisms for the coherent configuration $\mathcal{X}$} with respect to the algebraic isomorphism $\psi$.
Let ${\mathfrak{S}}\subseteq{\mathfrak{S}}_{m}$, where ${\mathfrak{S}}_{m}={\mathfrak{S}}_{m}({\mathcal{A}}^{\star})$. By formula in (48) and Lemma 8.3(2), there is a minimal (by inclusion) subset $\bar{V}=\bar{V}({\mathfrak{S}})$ of the set $V$ such that if ${\mathfrak{s}}\in{\mathfrak{S}}$, then ${\mathfrak{s}}\preceq{\mathfrak{s}}_{u,w}$ for some $u\in V_{1}\cap\bar{V}$ and $w\in V_{2}\cap\bar{V}$. Clearly,
@@
and hence $\Omega_{\bar{V}}\in F_{2m}({\mathcal{X}})$. Thus the $(2m)$-local system (51) contains an isomorphism $f_{\Delta}\in K_{V}$ of the coherent configuration $\mathcal{X}$}, where $\Delta=\Omega_{\bar{V}}$. Denote by ${\mathfrak{m}}({\mathfrak{S}})$ the inner $\mathcal{A}$}-multiplier $\mu(f_{\Delta})$.
@@
$\{{\mathfrak{m}}({\mathfrak{S}})\in{\mathfrak{M}}({\mathcal{A}}^{\star}):\ {\mathfrak{S}}\in{\mathfrak{S}}_{m}\}$
[…]
To verify the first relation in (28), let ${\mathfrak{S}}\in{\mathfrak{S}}_{m}$ and ${\mathfrak{s}}\in{\mathfrak{S}}$. By Lemma 8.3(2), we may assume without loss of generality that ${\mathfrak{s}}={\mathfrak{s}}_{u,w}$ for some vertices $u\in V_{1}\cap\bar{V}$ and $w\in V_{2}\cap\bar{V}$, where $\bar{V}=\bar{V}({\mathfrak{S}})$. Since $u,w\in\bar{V}$, we have $\Omega_{u,w}\subseteq\Omega_{\bar{V}}:=\Delta$. By the definition of $f_{\Delta}$, this implies that $(f_{\Delta})^{\Omega_{e}}\in\operatorname{Iso}({\mathcal{X}}_{\Omega_{e}},\psi_{\Omega_{e}})$, where $e=\{u,w\}$. By Lemma 9.4, either $e\not\in E$, or $e\in E$ and
@@
To verify the second relation in (28), let ${\mathfrak{S}},{\mathfrak{S}}^{\prime}\in{\mathfrak{S}}_{m}$. Put $\Delta=\Omega_{\bar{V}}$ and $\Delta^{\prime}=\Omega_{\bar{V}^{\prime}}$, where $\bar{V}=\bar{V}({\mathfrak{S}})$ and $\bar{V}^{\prime}=\bar{V}({\mathfrak{S}}^{\prime})$. As above, we have $\Delta,\Delta^{\prime}\in F_{2m}({\mathcal{X}})$. By the definition of $(2m)$-local system, this implies that
@@
for some $h_{\Delta,\Delta^{\prime}}\in\operatorname{Aut}({\mathcal{X}})$. It follows that $f_{\Delta}{f_{\Delta^{\prime}}}^{-1}=f_{\Delta,\Delta^{\prime}}h_{\Delta,\Delta^{\prime}}$, where the isomorphism $f_{\Delta,\Delta^{\prime}}:=f_{\Delta}{f_{\Delta^{\prime}}}^{-1}h_{\Delta,\Delta^{\prime}}^{-1}$ belongs to $K_{V}$ and is identical on $\Delta\cap\Delta^{\prime}$. Applying the monomorphism $\mu$ from Lemma 8.5 to the last equality, we obtain
@@
[…]